# Diving into a multi-band holographic superconductor

### v2

@@ 3.1 General structures:
For charged black hole with scalar hair, it can be shown that when the scalar hair forms, no matter how small it is, it can induce strong non-linear effect which removes the Cauchy horizonHartnoll:2020fhc ; Cai:2020wrp ; An:2021plu . There are many proofs of the no Cauchy horizon theorem by using conserved charge method and energy condition method Cai:2020wrp ; An:2021plu ; Yang:2021civ . We only briefly introduce the conserved charge method here. As the model has scaling symmetry, there is a conserved charge associated to this symmetry by Noether theorem. The conserved charge reads
$Q(z)=e^{\chi/2}[z^{-2}(fe^{-\chi})^{\prime}-\phi\phi^{\prime}].$ (20)
It can be directly verified that $Q^{\prime}(z)=0$ by using the equation of motion from Eq.(11) to Eq.(13). Assuming the black hole has two horizons at $z_{h}$ and $z_{i}$, thus the blackening factor $f(z)$ will satisfy the condition
$f(z_{h})=0,\quad f^{\prime}(z_{h})<0,$ (21)
$f(z_{i})=0,\quad f^{\prime}(z_{i})>0.$
There are also constraints on the gauge field by the regularity of equation of motion
$\phi(z_{h})=\phi(z_{i})=0$ (22)
By using the conditions (21) and (22), we can find
$Q(z_{h})=Q(z_{i})\to\frac{f^{\prime}(z_{h})}{z_{h}^{2}}e^{-\chi(z_{h})/2}=\frac{f^{\prime}(z_{i})}{z_{i}^{2}}e^{-\chi(z_{i})/2}$ (23)
This can not be true as the blackening factor must obey Eq.(21) as there are two horizons. Thus the assumption that the black hole has two horizons is wrong and there is only one horizon for black hole with scalar hair  [^4] This proof is also valid for $k=+1$ case, however we only focus on $k=0$ case in our paper as our holographic superconductor is in $k=0$ case. Note that this proof is valid for arbitrary numbers of minimally coupled scalar fields, thus there is no Cauchy horizon in our multi-band superconductor model and the black hole singularity becomes space-like when scalar charge forms. Absence of Cauchy horizon can also be corroborated by the numerical result, we plot the spacetime interior structure and gauge field behavior inside black hole in Fig.4 and Fig.5 to show this. It is easy to see that $h(z)=0$ and $\phi(z)=0$ have only one root which means the black hole has only one horizon.
Figure 4: Left panel: Metric function $h(z)$ inside black hole for the region not far from horizon. Right panel: Metric function $-h(z)$ inside black hole for the region far from horizon. We see $h(z)$ is an $O(1)$ quantity in the deep interior which means the singularity is space-like.
Figure 5: Gauge field $\phi(z)$ inside black hole. Left panel: not so far from horizon. Right panel: deep inside horizon.
3.2 ER collapse and Josephson oscillation:
Besides the absence of Cauchy horizon, there are also more novel behaviors inside the corresponding scalar hairy black hole. In this section, we will investigate the ER collapse and Josephson oscillation behavior. Near the horizon, there will be sudden decrease of metric function $g_{tt}$ which is so called ER collapse Hartnoll:2020rwq . This rapid collapse of $g_{tt}$ is a critical phenomenon which is only notable near the critical temperature $T_{c}$. We plot the ER collapse behavior in different temperature in Fig.6. It can be seen here that the ER collapse will be less dramatic as we lower the temperature away from $T_{c}$.
@@ 3.2 ER collapse and Josephson oscillation:
ER collapse is closely related to the phenomenon called Josephson oscillation. The scalar field inside black hole can have oscillations right after ER collapse happens. This can be found in terms of the equation of motion of scalar field  [^5] The two scalar fields have the same form of equation of motion,$\psi$ here denotes either $\psi_{1}$ or $\psi_{2}$ which reads
$\psi^{\prime\prime}+\frac{h^{\prime}}{h}\psi^{\prime}+\frac{\psi^{\prime}}{z}=-\frac{q^{2}\phi^{2}\psi}{z^{6}h^{2}}+\frac{m^{2}\psi}{z^{5}h}e^{-\chi/2}$ (24)
@@
After ER collapse, $e^{-\chi/2}$ is vastly small thus the mass term and $h^{\prime}/h$ term can be neglected. Moreover, it can be checked numerically in Fig.4 and Fig.5 that the $\phi^{2}/h^{2}$ term is approximately a constant after ER collapse. After using these approximations, the scalar field can be solved as
$\psi=c_{J}J_{0}(\frac{\phi}{2hz^{2}})+c_{Y}Y_{0}(\frac{\phi}{2hz^{2}})$ (25)
where $c_{J}$,$c_{Y}$ are two integration constants and $J_{0}$,$Y_{0}$ are first kind and second kind Bessel function respectively. Eq.(25) is a highly oscillating function which is denoted as "Josephson oscillation" in Ref.Hartnoll:2020fhc to show its connection to Josephson effect in superconductor.
@@
This analytical argument can also be corroborated by the numerical result. We plot the Josephson oscillation behavior of holographic multi-band superconductor in Fig.7 and Fig.8. For the temperature close to the critical temperature, we see that there is one scalar hair and as ER collapse is dramatic, the scalar field rapidly oscillates as shown in Fig.7.
Figure 7: Scalar oscillation behavior for $T=0.97T_{c}$, at this temperature the second scalar field $\psi_{2}$ has not appeared and the $\psi_{1}$ field rapidly oscillate.
As we lower the temperature, two things happen. Firstly, the ER collapse becomes less dramatic which leads to slower oscillation of scalar field. Secondly, below a certain temperature ($T=0.72360T_{c}$ in our case), second scalar field will appear. As we show in Fig.1 , for temperature $T=0.72T_{c}$, the temperature lies in the domain where there are two coexisted scalar fields. At this temperature, since the ER collapse is still not negligible, two scalar oscillations will appear as shown in the left panel of Fig.8. When further lowering the temperature, the ER collapse and Josephson oscillation cease to exist which is shown in the right panel of Fig.8.
Figure 8: Scalar oscillation behavior for $T=0.72T_{c}$ and $T=0.68T_{c}$. For $T=0.72T_{c}$ case, the $\psi_{1}$,$\psi_{2}$ coexist and oscillate. After further lowering the temperature, the oscillation disappears as a result of the lack of ER collapse.
3.3 Kasner epoch:
In the last section, we focus on the black hole interior structure near the horizon $r_{h}$. From the behavior of $z\psi^{\prime}$ in Fig. 9, we see that $z\psi^{\prime}_{1,2}$ approach constants in the deeper interior. This behavior will indicate that the system enters into the Kasner epoch,
@@ 3.3 Kasner epoch:
Figure 9: The scalar field behavior inside black hole. Left panel: we plot the behavior of $z\psi_{1}^{\prime}$. Right panel: we plot the behavior of $z\psi_{2}^{\prime}$. $z\psi^{\prime}$ approaches constant in the deep interior implies that the metric enters into the Kasner epoch.
@@
thus in this section we will turn to investigate the black hole interior structure in this deep region for multi-band case. We can check numerically that in this region, the mass term and charge term in the equations of motion can be neglected, and the equation of motion for $\psi_{1}$,$\psi_{2}$ and $\chi$ can be vastly simplified as follows
@@
After this approximation, the scalar field could be solved as
$\psi_{1}\sim\alpha_{1}\ln z,\quad\psi_{2}\sim\alpha_{2}\ln z$ (27)
for region near the singularity $z\to\infty$, where $\alpha_{1}$,$\alpha_{2}$ are two constants. The function $\chi$ can also be directly calculated as
$\chi\sim(\alpha_{1}^{2}+\alpha_{2}^{2})\ln z$ (28)
As can be verified from numerical results in Fig.4 posteriorly, $h=\frac{fe^{-\chi/2}}{z^{3}}$ is a $\mathcal{O}(1)$ quantity, thus we found
$f\sim z^{3+\frac{1}{2}\alpha_{1}^{2}+\frac{1}{2}\alpha_{2}^{2}}$ (29)
By using the metric ansatz in Eq.(7) the metric component $g_{zz}$ at large $z$ reads
$g_{zz}\sim\frac{1}{z^{5+\frac{1}{2}\alpha_{1}^{2}+\frac{1}{2}\alpha_{2}^{2}}}$ (30)
By transforming to proper time $\tau$ using the relation
$d\tau=\frac{dz}{z^{\frac{5}{2}+\frac{\alpha_{1}^{2}}{4}+\frac{\alpha_{2}^{2}}{4}}}$ (31)
we found that
$z\sim\tau^{-\frac{1}{\frac{3}{2}+\frac{1}{4}\alpha_{1}^{2}+\frac{1}{4}\alpha_{2}^{2}}}$ (32)
Thus the metric near the singularity reads
$ds^{2}=-d\tau^{2}+\tau^{2p_{t}}dt^{2}+\tau^{2p_{x}}dx^{2}+\tau^{2p_{y}}dy^{2}$ (33)
with Kasner exponent $p_{t}=\frac{\alpha_{1}^{2}+\alpha_{2}^{2}-2}{\alpha_{1}^{2}+\alpha_{2}^{2}+6}$ and $p_{x}=p_{y}=\frac{4}{\alpha_{1}^{2}+\alpha_{2}^{2}+6}$, it can be easily checked that Kasner exponents satisfy the relation
$p_{t}+p_{x}+p_{y}=1$ (34)
Also we can define $p_{\psi}$ as $\psi\sim-p_{\psi}\ln\tau$ and find that
$p_{\psi_{1}}=\frac{4\alpha_{1}}{\alpha_{1}^{2}+\alpha_{2}^{2}+6},\quad p_{\psi_{2}}=\frac{4\alpha_{2}}{\alpha_{1}^{2}+\alpha_{2}^{2}+6}$ (35)
Thus we find another relation among Kasner exponents,
$p_{t}^{2}+p_{x}^{2}+p_{y}^{2}+p_{\psi_{1}}^{2}+p_{\psi_{2}}^{2}=1.$ (36)
@@
$\phi\sim\phi_{0}+E_{0}z^{1-\frac{1}{2}\alpha_{1}^{2}-\frac{1}{2}\alpha_{2}^{2}}$ (37)
@@
In the Kasner region, as verified posterity that the term involving $e_{1}$ and $e_{2}$ and $m$ in all the equations of motion can be neglected. The Maxwell equation and Einstein equation involving $h$ can be solved as
[…]
@@
$\chi^{\prime}=\frac{\alpha_{1}(z)^{2}+\alpha_{2}(z)^{2}}{z}$ (41)
Plugging Eq.(39-41) into the scalar field equations, the scalar field equations become
$\alpha_{2}^{2}+\alpha_{1}^{2}+2z\frac{\alpha_{1}^{\prime\prime}}{\alpha_{1}^{\prime}}-4z\frac{\alpha_{1}^{\prime}}{\alpha_{1}}=0$ (42)
$\alpha_{1}^{2}+\alpha_{2}^{2}+2z\frac{\alpha_{2}^{\prime\prime}}{\alpha_{2}^{\prime}}-4z\frac{\alpha_{2}^{\prime}}{\alpha_{2}}=0$ (43)
Note that when one of the scalar fields vanishes, the above differential equations becomes one single equation
$\alpha^{2}+2z\frac{\alpha^{\prime\prime}}{\alpha^{\prime}}-4z\frac{\alpha^{\prime}}{\alpha}=0$ (44)
@@
$(\alpha-\alpha_{0})^{-2/(2-\alpha_{0}^{2})}(\frac{2}{\alpha_{0}}-\alpha)^{-1/(1-2/\alpha_{0}^{2})}\alpha=\frac{z_{in}}{z}$ (45)
where $z_{in}$ is an integration constant. This solution tells us that if the first Kasner epoch has the feature $\alpha_{0}<\sqrt{2}$, when increasing $z$, $\alpha(z)$ should transit from $\alpha_{0}$ to $2/\alpha_{0}$. This is the so called "Kasner inversion" behavior which is first introduced in Ref. Hartnoll:2020fhc . [^6] The Kasner transition behaviors are different for different matter fields or gravitational theories, for more discussions see Ref.An:2022lvo ; Cai:2024ltu ; Caceres:2024edr ; Bueno:2024fzg . However, the presence of second scalar field will change this inversion behavior, the second scalar field will make the differential equation become the coupled differential equations. The differential equations Eq.(42)and Eq.(43) may not be directly solved analytically, but in all the cases where Kasner transition happens, we can find numerically that the Kasner transition has two interesting relations
$\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}=2,\quad\alpha_{1}\beta_{2}=\alpha_{2}\beta_{1}$ (46)
@@
where $\alpha_{1}$ and $\beta_{1}$ are the values of $z\psi^{\prime}_{1}$ before and after transition, and $\alpha_{2}$ and $\beta_{2}$ are the values of $z\psi_{2}^{\prime}$ before and after transition. We show one example of this Kasner transition behavior in Fig.10, and more cases are presented in appendix A in Fig.12. The results are also summarized in the Table 1 .
@@
It is interesting to check that these two conditions are crucial to make the Kasner transition follow the BKL transition rule Belinsky:1970ew ; Damour:2002et which is
$p_{t}\to-\frac{p_{t}}{2p_{t}+1},\quad p_{x}\to\frac{p_{x}+2p_{t}}{2p_{t}+1},\quad p_{y}\to\frac{p_{y}+2p_{t}}{2p_{t}+1}$ (47)
We see that $p_{t}$ changes signs during the transition period. Before the transition, $p_{t}$ is negative as a result of $\alpha_{1}^{2}+\alpha_{2}^{2}<2$ while $p_{t}$ is positive after the transition. Therefore, $p_{t}<0$ which physically means that spatial $t$ direction of the geometry expands is an unstable period while $p_{t}>0$ is a stable period. As proved in Ref.Henneaux:2022ijt , as long as scalar fields are minimally coupled, there is final Kasner regime no matter how many scalar fields are. So there will be no more Kasner transition after entering into $p_{t}>0$ region.
@@
3.4 Thermal a-function:
The difference between Kasner exponents we found in Fig.11 can also be reflected in other physical quantities. In this section, we choose the thermal a-function which was originally raised in Caceres:2022smh to illustrate this.
For AdS black hole, the radial coordinate $z$ is commonly treated as an energy scale which parameterized the renormalization group flow from the UV CFT on the boundary to the IR CFT on the horizon. It was shown in Caceres:2022smh that by analytically continuation, the holographic renormalization group flow can also be generalized to describe the black hole interior region, which is called trans-IR flow. Along the flow from UV to IR and then to trans-IR region,a monotonic function can be defined which is called thermal a-function in Caceres:2022smh . The thermal a-function measures the effective number of degrees of freedom for each energy scale along the RG flow whose expression in $d+1$ dimensions is
$a_{T}(u)=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2})\kappa^{d-1}}(\frac{\sqrt{f(u)}e^{-\chi/2}}{A^{\prime}(u)})^{d-1}$ (48)

### v3

@@ 3.1 General structures:
For charged black hole with scalar hair, it can be shown that when the scalar hair forms, no matter how small it is, it can induce strong non-linear effect which removes the Cauchy horizon[23, 22, 25]. There are many proofs of the no Cauchy horizon theorem by using conserved charge method and energy condition method [22, 25, 43]. The proof is valid for arbitrary numbers of minimally coupled scalar fields, thus there is no Cauchy horizon in our multi-band superconductor model and the black hole singularity becomes space-like when scalar condensate forms. Absence of Cauchy horizon can also be corroborated by the numerical result, we plot the spacetime interior structure and gauge field behavior inside black hole in Fig.3 to show this. It is easy to see that $h(z)=0$ and $\phi(z)=0$ have only one root which means the black hole has only one horizon.
(a) Plot of metric function $-h(z)$
(b) Plot of gauge field $\phi(z)$
Figure 3: We see $h(z)$ is an $O(1)$ quantity in the deep interior which means the singularity is space-like.
3.2 Interior structure near the horizon:
Besides the absence of Cauchy horizon, there are also more novel behaviors inside the corresponding scalar hairy black hole. Notable effects include the ER collapse and Josephson oscillation of scalar field [23]. Thus we will first investigate them for multi-band model.
Near the horizon, similar to the results in Ref.[23], there will still be sudden decrease of metric function $g_{tt}$ which is called ER collapse. We plot the ER collapse behavior in different temperature in Fig.4. It can be seen that the ER collapse will be less dramatic as we decrease the temperature away from $T_{c}$.This means that the collapse of $g_{tt}$ is a critical phenomenon which is only notable near the critical temperature $T_{c}$.
@@ 3.2 Interior structure near the horizon:
ER collapse is closely related to the phenomenon called Josephson oscillation. The scalar field inside black hole can have oscillations right after ER collapse happens. This can be found in terms of the equation of motion of scalar field  [^5] The two scalar fields have the same form of equation of motion,$\psi$ here denotes either $\psi_{1}$ or $\psi_{2}$. which reads
$\psi^{\prime\prime}+\frac{h^{\prime}}{h}\psi^{\prime}+\frac{\psi^{\prime}}{z}=-\frac{q^{2}\phi^{2}\psi}{z^{6}h^{2}}+\frac{m^{2}\psi}{z^{5}h}e^{-\chi/2}.$ (16)
@@
After ER collapse, $e^{-\chi/2}$ is vastly small thus the mass term and $h^{\prime}/h$ term can be neglected. Moreover, it can be checked numerically in Fig.3 that the $\phi^{2}/h^{2}$ term is approximately a constant after ER collapse. After using these approximations, the scalar field can be solved as
$\psi=c_{J}J_{0}(\frac{\phi}{2hz^{2}})+c_{Y}Y_{0}(\frac{\phi}{2hz^{2}}),$ (17)
@@
where $c_{J}$,$c_{Y}$ are two integration constants and $J_{0}$,$Y_{0}$ are first kind and second kind Bessel function respectively. Eq.(17) is a highly oscillating function which is called ”Josephson oscillation” in Ref.[23] to show its close connection to Josephson effect in superconductor theory. This analytical argument can also be corroborated by the numerical result. We plot the Josephson oscillation behavior of holographic multi-band superconductor in Fig.5. For the temperature close to the critical temperature, we see that there is one scalar hair which oscillates rapidly. As we lower the temperature, two things happen. Firstly, the ER collapse becomes less dramatic which leads to slower oscillation of scalar field. Secondly, below a certain temperature ($T=0.72360T_{c}$ in our case), second scalar field appears. As was shown in the middle panel of Fig.5, for temperature $T=0.72T_{c}$, there will be two oscillating scalar fields. Further decreasing the temperature will eliminate the Josephson oscillation which is shown in the right panel of Fig.5.
From the above results, we conclude that the interior structure of holographic multi-band superconductor near the horizon is quite similar to the results for single-band case. However, we will show in the following that the physical properties near the singularity will be quite different.
(a) Scalar oscillation behavior for $T=0.97T_{c}$
(b) Scalar oscillation behavior for $T=0.72T_{c}$
(c) Scalar oscillation behavior for $T=0.68T_{c}$
Figure 5: Josephson oscillation for different temperature. When lowering temperature, the oscillation gradually disappears as ER collapse gradually diminishes.
3.3 Interior structure near the singularity:
From the behavior of $z\psi^{\prime}$ in Fig.6, we see that the scalar field will change logarithmically with respect to $z$ (which is equivalent to $z\psi^{\prime}_{1,2}$ approach constants) in the deep interior. This behavior indicates that the system enters into the new epoch.
(a) Behavior of $z\psi_{1}^{\prime}$
(b) Behavior of $z\psi_{2}^{\prime}$
@@ 3.3 Interior structure near the singularity:
Figure 6: The scalar field behavior inside black hole. $z\psi^{\prime}$ approaches constant in the deep interior implies that the metric enters into the Kasner epoch.
@@
We can check numerically that in this region, the mass term and charge term in the equations of motion can be neglected, and the equation of motion for $\psi_{1}$,$\psi_{2}$ and $\chi$ can be vastly simplified as follows
@@
After this approximation, the scalar fields and the function $\chi$ can be solved as
$\psi_{1}\sim\alpha_{1}\ln z,\quad\psi_{2}\sim\alpha_{2}\ln z,\quad\chi\sim(\alpha_{1}^{2}+\alpha_{2}^{2})\ln z$ (19)
where $\alpha_{1}$,$\alpha_{2}$ are two integration constants. As can be verified from numerical results in Fig.3 posteriorly, $h=\frac{fe^{-\chi/2}}{z^{3}}$ is a $\mathcal{O}(1)$ quantity, thus we find the metric function $f$ will have following form
$f\sim z^{3+\frac{1}{2}\alpha_{1}^{2}+\frac{1}{2}\alpha_{2}^{2}}.$ (20)
By using the metric ansatz in Eq.(3) , the metric component $g_{zz}$ at large $z$ reads
$g_{zz}\sim\frac{1}{z^{5+\frac{1}{2}\alpha_{1}^{2}+\frac{1}{2}\alpha_{2}^{2}}}.$ (21)
By transforming to proper time $\tau$ using the relation $d\tau=\sqrt{g_{zz}}dz$, we found that the metric near the singularity reads
$ds^{2}=-d\tau^{2}+\tau^{2p_{t}}dt^{2}+\tau^{2p_{x}}dx^{2}+\tau^{2p_{y}}dy^{2},$ (22)
with Kasner exponent
$p_{t}=\frac{\alpha_{1}^{2}+\alpha_{2}^{2}-2}{\alpha_{1}^{2}+\alpha_{2}^{2}+6},\quad p_{x}=p_{y}=\frac{4}{\alpha_{1}^{2}+\alpha_{2}^{2}+6}.$ (23)
Moreover, we can define $p_{\psi}$ as $\psi\sim-p_{\psi}\ln\tau$ and find that
$p_{\psi_{1}}=\frac{4\alpha_{1}}{\alpha_{1}^{2}+\alpha_{2}^{2}+6},\quad p_{\psi_{2}}=\frac{4\alpha_{2}}{\alpha_{1}^{2}+\alpha_{2}^{2}+6}.$ (24)
Thus with two independent scalar fields, the Kasner exponents should be determined jointly by the two parameters $\alpha_{1}$ and $\alpha_{2}$. It can be easily checked that the above Kasner exponents satisfy the following two interesting relations
$p_{t}+p_{x}+p_{y}=1,\quad p_{t}^{2}+p_{x}^{2}+p_{y}^{2}+p_{\psi_{1}}^{2}+p_{\psi_{2}}^{2}=1.$ (25)
@@
$\phi\sim\phi_{0}+E_{0}z^{1-\frac{1}{2}\alpha_{1}^{2}-\frac{1}{2}\alpha_{2}^{2}},$ (26)
@@
In the Kasner region, as verified posterity that the terms involving $e_{1}$ and $e_{2}$ and $m$ in all the equations of motion can be neglected, the Maxwell equation (7) and Einstein equation (9) can be solved as
@@
$\chi^{\prime}=\frac{\alpha_{1}(z)^{2}+\alpha_{2}(z)^{2}}{z}.$ (29)
Plugging Eq.(27-29) into the scalar field equations, the scalar field equations become
$\alpha_{2}^{2}+\alpha_{1}^{2}+2z\frac{\alpha_{1}^{\prime\prime}}{\alpha_{1}^{\prime}}-4z\frac{\alpha_{1}^{\prime}}{\alpha_{1}}=0,$ (30)
$\alpha_{1}^{2}+\alpha_{2}^{2}+2z\frac{\alpha_{2}^{\prime\prime}}{\alpha_{2}^{\prime}}-4z\frac{\alpha_{2}^{\prime}}{\alpha_{2}}=0.$ (31)
Note that when one of the scalar fields vanishes, the above coupled differential equations becomes one single differential equation
$\alpha^{2}+2z\frac{\alpha^{\prime\prime}}{\alpha^{\prime}}-4z\frac{\alpha^{\prime}}{\alpha}=0,$ (32)
@@
$(\alpha-\alpha_{0})^{-2/(2-\alpha_{0}^{2})}(\frac{2}{\alpha_{0}}-\alpha)^{-1/(1-2/\alpha_{0}^{2})}\alpha=\frac{z_{in}}{z},$ (33)
where $z_{in}$ is an integration constant. This solution tells us that if the first Kasner epoch has the feature $\alpha_{0}^{2}<2$, when increasing $z$, $\alpha(z)$ should transit from $\alpha_{0}$ to $2/\alpha_{0}$. This is the so called ”Kasner inversion” behavior which is first introduced in Ref.[23]. [^6] The Kasner transition behaviors are different for different matter fields or gravitational theories, for more discussions see Ref.[28, 31, 39, 44]. However, for the multi-band model in the presence of second scalar field, the simple Kasner inversion behavior will change significantly. In the multi-band case, the two scalar fields are both crucial to determine the transition behavior which is reflected in the coupled differential equations Eq.(30)and Eq.(31). The differential equations may not be directly solved analytically, but in all the cases where Kasner transition happens, we can find numerically that the Kasner transition has two interesting relations
$\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}=2,\quad\alpha_{1}\beta_{2}=\alpha_{2}\beta_{1}.$ (34)
@@
where $\alpha_{1}$ and $\beta_{1}$ are the values of $z\psi^{\prime}_{1}$ before and after transition, and $\alpha_{2}$ and $\beta_{2}$ are the values of $z\psi_{2}^{\prime}$ before and after transition. We show one example of this Kasner transition behavior in Fig.7, and more cases are presented in A in Fig.9. The results are also summarized in the Table 1. The relations we find in Eq.(34) are the first Kasner transition rule in the presence of two free parameters. It should be distinguished from other Kasner transition rules in Ref.[28, 31, 34] where there is only one free parameter.
@@
It is interesting to check that these two conditions are crucial to make the Kasner transition follow the BKL transition rule [40, 41, 42] which is
$p_{t}\to-\frac{p_{t}}{2p_{t}+1},\quad p_{x}\to\frac{p_{x}+2p_{t}}{2p_{t}+1},\quad p_{y}\to\frac{p_{y}+2p_{t}}{2p_{t}+1}.$ (35)
We see that $p_{t}$ changes signs during the transition period. Before the transition, $p_{t}$ is negative as a result of $\alpha_{1}^{2}+\alpha_{2}^{2}<2$ while $p_{t}$ is positive after the transition. Therefore, $p_{t}<0$ which physically means that spatial $t$ direction of the geometry expands is an unstable period while $p_{t}>0$ is a stable period. As proved in Ref.[45], as long as scalar fields are minimally coupled, there is final Kasner regime no matter how many scalar fields are. So there will be no more Kasner transition after entering into $p_{t}>0$ region.
@@
4 Conclusions and Discussions
In this work, we investigate the interior structure of holographic multi-band superconductor. For the region near the horizon, there still exists ER collapse behavior after which Josephson oscillation occurs. When lowering the temperature below which second scalar condensate appears, there will be oscillation behavior of two scalar fields. For the region near the singularity, the additional scalar field has contributions to the Kasner exponent which makes the Kasner exponent different from the single scalar case. In this multi-band case, the Kasner exponents are determined by two independent parameters $\alpha_{1}$ and $\alpha_{2}$ which is more complex than the single-band case. This feature affects the Kasner transition behavior significantly, which makes the Kasner transition present distinct relations compared to the single band case. The transition rule we found in Eq.(34) is the first kind of Kasner transition in the presence of two independent parameters which will be crucial for characterizing the general interior structure of hairy black holes. Furthermore, these results certainly illuminate that the black hole interior plays crucial role in describing the boundary condensed matter system.
In this work, we only focus on the s-wave superconductor. However, p-wave superconductor can have more intricate interior structure compared to s-wave case [27]. So it will be interesting to investigate the interior of multi-band model with coexistence of s-wave and p-wave order parameters [16, 15] which may present more novel phenomena.
Moreover, it is interesting to ask if the near singularity structure we found in Fig.8 can be directly reflected in some physical quantities of boundary systems. Based on the proposal raised in [46], we can claim that the structure we found can at least be closely related to the renormalization group theory of boundary system.
For AdS black hole, the radial coordinate $z$ is commonly treated as an energy scale which parameterized the renormalization group flow from the UV CFT on the boundary to the IR CFT on the horizon. It was shown in [46] that by analytically continuation, the holographic renormalization group flow can also be generalized to describe the black hole interior region, which is called trans-IR flow. Along the flow from UV to IR and then to trans-IR region, a monotonic function can be defined which is called thermal a-function [46]. The thermal a-function measures the effective degrees of freedom of field theory at each energy scale along the RG flow whose expression in $d+1$ dimensions is
$a_{T}(u)=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2})\kappa^{d-1}}(\frac{\sqrt{f(u)}e^{-\chi/2}}{A^{\prime}(u)})^{d-1},$ (36)